A jeep leaves Orlando traveling 40 mph. 20 minutes later a car leaves Orlando following the jeep traveling 55 mph. How long will it take, in minutes for the car to overtake or pass the jeep?
Answer by lwsshak3(11628)

A jeep leaves Orlando traveling 40 mph. 20 minutes later a car leaves Orlando following the jeep traveling 55 mph. How long will it take, in minutes for the car to overtake or pass the jeep?
***
let x=travel time of car
x+20 min=x+1/3 hr=travel time of jeep
speed*travel time=distance(same for both jeep and car)
..
40(x+1/3)=55x
40x+40/3=55x
lcd:3
120x+40=165x
45x=40
x=40/45=8/9 hr=53 1/3 min
How long will it take for the car to overtake or pass the jeep? 53 1/3 minutes